OpenStudy (anonymous):

Title: osmosis Aim: To calculate the solute potential of plant cells. Q.: How do I calculate the solute potential of each onion epidermal cell listed in the table below?

OpenStudy (anonymous):



OpenStudy (anonymous):

I was given a table of fixed values of the 'water potential' and the respective conc. of NaCl...I don't understand how those values were obtained but the thing is, the conc. starts at 0.1 mol/dm^3 and ends at 1.0 mol/dm^3. However, when my classmates and I did the lab, we had to prepare NaCl solutions of different concentrations. We ended up having concs. of 5.7, 4.6, 2.9, 3.4 and 5.7. How do I calculate the solute potential?

OpenStudy (frostbite):

I suggest you use this: \[\Large \Delta G=RT \ln \left( \frac{ C_i }{ C_o } \right)\] Where \(C_{i}\) is the concentration inside the cell, and \(C_{o}\) is the concentration out side the cell. in equilibrium \(\Delta G=0\) obviously. If you want the osmotic pressure, we can use the van't Hoff equation and assume ideal solution and write: \[\Large \Pi=[J]RT\] Where is \([j]\) is the concentration of the active specie, for Nacl we then get \(2[j]\) as Na and Cl will add equality to the osmotic pressure. Sorry I'm not more of a help, but I am not known with the term "solute potential".

OpenStudy (anonymous):

I understand. How does one know the concentration inside a cell?

OpenStudy (nincompoop):

\[ \large \psi_{water~potential} = \psi_{pressure~potential}+\psi_{solute~potential} \\ \large \color{red}{ \rightarrow \psi_{sp} = \psi_{wp}-\psi_{pp}} \\ where: \psi_{sp} = iCRT \\ \]

OpenStudy (nincompoop):

http://5e.plantphys.net/article.php?id=29